An urn contains 10 white and 6 black balls. Balls are randomly selected, one at a time, until a black one is obtained. If we assume that each ball selected is replaced before the next one is drawn, what is the probability that

a) exactly 5 draws are needed?

b) at least 3 draws are needed?
Answer #1

P(black) = 6/16 = 3/8

P(white) = 10/16 = 5/8

a) P(exactly 5 draws are needed) = P(first 4 draws white) x P(5th draw is black)

= (5/8)4 x 3/8

= 0.0572

b) P(at least 3 draws are needed) = P(first 2 draws are white)

= (5/8)2

= 0.3906